OpenStudy (anonymous):

Determine the delivery radius for your shop. Draw a point on a coordinate plane where your shop will be located. Create two different radii lengths from your shop, and construct the circles that represent each delivery area.

OpenStudy (anonymous):

a) How much area will each delivery radius cover? b) Write the equation for each circle created.

OpenStudy (amistre64):



OpenStudy (amistre64):

determine a center and a radius and write up the equations: center (a,b), radius r \[(x-a)^2 + (y-b)^2 = r^2\]

OpenStudy (anonymous):

how do I do that again @amistre64

OpenStudy (amistre64):

i thought my posts were pretty well self explantory

OpenStudy (anonymous):

Okay, but I'm still confused. How do I find the area?

jimthompson5910 (jim_thompson5910):

to find the circle's area, use this formula A = pi*r^2

OpenStudy (anonymous):

okay. With what numbers?

OpenStudy (amistre64):

this question is asking YOU to pick numbers ... not us.

OpenStudy (amistre64):

you need to determine you favorite center point, and your favorite 2 radius lengths ... the rest is just plug and play

OpenStudy (anonymous):

I KNOW that much. But HOW do I figure out WHERE to plug them in?

OpenStudy (amistre64):

determine a center and a radius and write up the equations: center (a,b), radius r \[ (x-a)^2 + (y-b)^2 = r^2\]

OpenStudy (anonymous):

WHY do you keep typing that equation in!!! I already know that and I asked @jim_thompson5910 to help me. Because your confusing ME!!!

OpenStudy (amistre64):

... good luck then

jimthompson5910 (jim_thompson5910):

well if you wanted to find the area of a circle with a radius of 6, then you would replace r with 6 in A = pi*r^2 but you can come up with circles of any radius since it says "Create two different radii lengths"

OpenStudy (anonymous):

So, if I had the coordinates (3,6), then how would I plug that in?

jimthompson5910 (jim_thompson5910):

for the center you mean?

OpenStudy (anonymous):

YEA.

jimthompson5910 (jim_thompson5910):

if so, then (3,6) = (a,b) as amistre64 is saying so a = 3 b = 6 and you plug them into (x-a)^2 + (y-b)^2 = r^2

OpenStudy (anonymous):

By the way, @amistre64 I didn't mean it in a rude way. I just need to finish the assignment because I have a deadline and I didn't understand what you were saying Don't take it offensive

OpenStudy (anonymous):

And for the radius, how would I do that. Find the diameter and divide it in half @jim_thompson5910

jimthompson5910 (jim_thompson5910):

the radius is something you come up with since it says "Create two different radii lengths..."

jimthompson5910 (jim_thompson5910):

so this is where you can be creative

OpenStudy (anonymous):

OHHH.... That's why @amistre64 kept saying "Create two different radii lengths?" So. I could choose 7, and that would be my radius

jimthompson5910 (jim_thompson5910):

yep, any positive number you want

OpenStudy (anonymous):

thank you. I have to create a second circle

jimthompson5910 (jim_thompson5910):

np

OpenStudy (anonymous):

wait @jim_thompson5910 I was saying. am I supposed to create a second circle?

jimthompson5910 (jim_thompson5910):

oh yes, yes you do

OpenStudy (amistre64):

im fine :)

OpenStudy (anonymous):

just do the same thing, as before @jim_thompson5910 ?

jimthompson5910 (jim_thompson5910):

I'm assuming it's from the same center, so just pick a different radius

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

@amistre64 , okay, because I thought I had offended you and I didn't mean to because I wasn't understanding what you was saying. I do now because it was explained to me. thanks for putting up with my confusion :) :)

OpenStudy (anonymous):

thanks, @jim_thompson5910 enjoy ur medal

jimthompson5910 (jim_thompson5910):

yw